OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

I assume you mean (3/2)x here. So the slope is 3/2. we plug that and the desired point into y=mx+b 3=(3/2)(-6)+b and solve for b to find b=12, so y=(3/2)x +12